A 30 degree angle cuts off an arc in a circle with a diameter of 10 in. find the length of the arc to the nearest hundreth.
1
Expert's answer
2011-08-31T06:25:54-0400
The arc length is S= Rθ , where θ is the angle measured in radians.
In our case θ = 30 degree = π/6
R = 10/2 = 5 in.
Thus
S = 5π/6 = 2.62 in.